MOLECULAR CONSTANTS OF C_{2} IN THE c^{3}Σ_{u}^{+} STATE

\underline{MASAKAZU NAKAJIMA} and YASUKI ENDO, Department of Basic Science, The University of Tokyo, Tokyo 153-8902, Japan.

The d^{3}Π_{g} − c^{3}Σ_{u}^{+} band system of C_{2} was observed in a discharge flow cell by laser induced fluorescence spectroscopy. Although the observation of the band system has been reported for the jet-cooled C_{2} molecule, ^{a} transitions involving higher rotational levels were observed in the present study. Improved molecular constants for several vibrational levels of the c^{3}Σ_{u}^{+} state will be presented. Furthermore, the bands with v′′ = 5, 6, 7 were newly observed in a discharged supersonic jet, of which molecular constants were determined for the first time.

aJ. A. Joester et al., J. Chem. Phys. 127, 214303 (2007).